# Improving LLM First-Token Predictions in Multiple-Choice Question Answering via Prefilling Attack

### v1

@@ 4 Experiments
MMLU RACE OBQA SIQA MS CQA SC HS MC-T SciQ MQA ARC-E LQA $\bar{\Delta}$
Llama-3.1-8B 63.1 78.1 71.2 70.9 87.0 72.7 93.2 52.2 82.1 97.1 24.0 90.2 28.9
+ prefilling 68.4 83.2 84.8 72.3 93.2 76.5 96.4 69.0 82.5 98.3 33.8 94.7 34.0 +5.9
Qwen-2-7B 68.9 87.0 83.8 75.6 88.1 79.4 97.1 53.4 88.3 97.5 29.8 94.3 33.0
+ prefilling 69.1 86.7 84.8 76.1 92.6 79.4 97.3 65.2 86.1 97.4 36.5 94.7 35.5 +1.9
Gemma-7B 45.9 50.4 40.6 34.2 77.5 51.6 69.6 36.3 60.1 89.0 24.1 55.9 26.4
      + prefilling 52.4 70.5 71.2 65.2 64.3 68.6 92.0 56.3 91.0 95.1 25.1 86.7 26.1 +15.6
Gemma-2-9B 33.9 49.5 38.8 56.3 85.9 39.6 97.6 43.0 72.2 31.8 27.8 38.6 31.0
      + prefilling 72.1 86.3 89.8 74.2 87.3 79.0 97.9 69.8 78.5 98.3 34.0 96.6 35.8 +27.2
Zephyr-7B 57.5 86.5 73.0 67.5 83.6 66.0 96.0 35.0 69.1 88.5 23.3 95.6 32.1
      + prefilling 58.8 87.3 70.8 68.5 86.2 69.0 96.0 36.1 93.3 93.1 24.3 98.1 40.6 +3.7
Ministral-8B 62.1 84.7 82.4 74.7 87.3 74.3 97.4 81.1 91.5 97.4 23.1 93.6 28.4
      + prefilling 63.9 84.7 85.2 75.9 89.9 75.0 98.2 86.5 91.9 98.0 24.5 93.7 34.6 +1.9
Mistral-Nemo-12B 65.1 82.7 80.6 74.2 78.0 74.1 97.3 51.6 92.4 96.8 23.5 92.6 28.7
      + prefilling 66.0 83.1 80.8 75.8 80.9 75.4 96.9 76.4 93.3 97.2 25.4 93.1 32.0 +3.0
Phi-4-14B 76.4 73.3 84.0 71.7 83.4 72.5 98.2 61.2 82.5 95.4 25.0 87.8 31.9
      + prefilling 79.7 73.4 90.0 77.3 93.2 79.6 98.7 87.8 86.1 98.3 45.0 87.8 34.6 +6.8
Table 1: Benchmarking the fundamental abilities of base language models on natural language understanding, general question answering, coding, mathematics, scientific knowledge, reasoning, etc. Prefilling each model response positively impacts its ability to output the correct option.
4.1 Experimental Setup
Benchmarks. We evaluate our approach across several widely recognized MCQA benchmarks, each designed to test different cognitive capabilities and domains of knowledge. We begin with general knowledge, assessed using MMLU (Hendrycks et al., 2021), a comprehensive benchmark covering 57 academic and professional subjects. For reading comprehension, we employ RACE (Lai et al., 2017), a dataset derived from middle and high school exams and OpenBookQA (OBQA) (Mihaylov et al., 2018), which requires reasoning over scientific facts and general knowledge. In the domain of commonsense reasoning, we include Social IQa (SIQA) (Sap et al., 2019) for social interaction understanding, Moral Stories (MS) (Emelin et al., 2020) for moral decision-making and CommonsenseQA (CQA) (Talmor et al., 2019) for everyday commonsense reasoning. For narrative understanding, we consider Story Cloze (SC) (Mostafazadeh et al., 2016), where the model selects the most plausible story continuation; HellaSwag (HS) (Zellers et al., 2019), which evaluates narrative coherence in the presence of adversarial distractors, and MC-TACO (MC-T) (Zhou et al., 2019), which assesses temporal and causal reasoning in short texts. In the area of STEM, we use SciQ (Welbl et al., 2017) for scientific knowledge and MathQA (MQA) (Amini et al., 2019) to evaluate mathematical problem-solving. For logical and analytical reasoning, we employ ARC-Easy (ARC-E) (Clark et al., 2018), which focuses on elementary-level scientific reasoning and recall, and LogiQA (LQA) (Liu et al., 2020), based on national civil service exams. Together, these datasets enable a comprehensive evaluation of our method effectiveness and robustness across varied cognitive demands.
Models. We assess a range of open-source instruction-tuned language models, selected to cover diverse architectures, sizes, and alignment strategies. Our pool includes Meta’s Llama-3.1-8B, Alibaba’s Qwen-2-7B, DeepMind’s Gemma-7B and Gemma-2-9B, HuggingFace’s DPO-tuned Zephyr-7B, Microsoft’s efficient Phi-4-14B, and two Mistral-family models: Ministral-8B and the alignment-focused Mistral-Nemo-12B. This selection enables a comprehensive analysis of prefilling effectiveness across instruction-tuned models.
Prompt Format. For each benchmark, inputs are formatted as a concatenation of general_instruction, question, and answers. To apply our prefilling strategy, we insert a natural-language prefix at the beginning of the model response. The default template used is: ‘Given the question and the possible options, my answer is: ’ that, placed immediately after the assistant message tag in chat-based formats, promotes symbolic alignment and encourages structured, interpretable outputs. This template was selected based on preliminary trials, as it consistently yielded strong performance across all benchmarks and proved robust across different models. We consider as valid answers only those token sequences that match the symbolic answer labels exactly (e.g., ‘A’, ‘B’, ‘C’, ‘D’), optionally preceded by up to two whitespace or newline characters. A thorough analysis on the robustness of the proposed strategy to different templates are provided in the Appendix (cf. Table 6).
@@ 4.1 Experimental Setup
We begin by evaluating the overall effectiveness of our prefilling strategy in enhancing the answer accuracy of the model across a diverse set of MCQA benchmarks, by comparing it with a standard FTP approach. Specifically, for standard FTP, the model is prompted with the task and options, and the next-token probability distribution is used to score the first-token likelihoods assigned to each answer option. No prefix is injected. When using our prefilling strategy, instead, a fixed natural-language prefix is injected at the beginning of each assistant response before the model generates any output. The FTP scoring is then applied exactly as in the baseline. This approach retains the efficiency of FTP but adds a steering mechanism that helps the model align with the expected symbolic format.
Results. Table 1 reports first-token accuracy for each model across benchmarks, comparing performance with and without prefilling. For the setting with prefilling, we always use the default prefilling template. As it can be seen, our prefilling strategy leads to consistent improvements across all models and datasets. In particular, the largest gain is achieved by Gemma-2-9B, which improves by over +27 points on average. Notably, Gemma-7B and Zephyr-7B improve by +15.6 and +3.7 points, respectively. Even strong baseline models like LLama-3.1-8B and Phi-4-14B show meaningful improvements (i.e., +5.9 and +6.8, respectively), demonstrating that prefilling benefits models regardless of their initial performance level.
These results confirm that structured prefilling robustly improves alignment between the model’s most probable output and the intended symbolic answer, reducing both misalignment and misinterpretation errors.
@@ 4.3 Alignment to Open-Ended Generation
To further contextualize our results, we compare our approach against an open-ended generation setting, which we consider a gold-standard reference due to its flexibility and alignment with natural model behavior – albeit at higher computational, training, maintenance, and evaluation cost due to decoding and external LLM classifier needs. In this setting, models generate free-form answers to each question without being constrained to select from symbolic options.
These open-ended responses are then mapped to symbolic answer labels through automatic evaluation using language model classifiers, such as GPT-3.5-Turbo (Ouyang et al., 2022) and Llama-3.1-70B-Instruct (Dubey et al., 2024). The aim of this comparison is to measure how well our prefilling strategy approximates the answers that a model would naturally produce in an unconstrained, expressive scenario. As in previous settings, we use the accuracy metric to quantify model performance by evaluating whether the extracted symbolic answer matches the ground truth.
Results. Figure 3 summarizes results on three representative benchmarks: OpenBookQA, Social IQa, and SciQ [^1] The complete results on all benchmarks are reported in Section A.2, as well as the exact prompting and classification details for open-ended responses.. We observe that FTP with our prefilling strategy consistently achieves higher alignment with the symbolic labels derived from open-ended generation, outperforming the standard FTP baseline across all tasks. In several cases, FTP with prefilling even surpasses the accuracy of open-ended generation itself as judged by GPT-based evaluation, suggesting that it provides not only a more accurate but also a more efficient surrogate for expressive generation.
These results reinforce the idea that prefilling bridges the gap between efficient token-level scoring and natural free-form output, enabling structured symbolic accuracy without sacrificing alignment to the model’s unconstrained reasoning.
Figure 3: Comparison of model accuracy on OpenBookQA, Social IQa, and SciQ using standard FTP, FTP with prefilling, and open-ended generation with GPT-3.5-Turbo or Llama-3.1-70B as classifiers. FTP with prefilling consistently outperforms baseline FTP and often surpasses the more expensive open-ended generation approaches.
@@ 4.4 Analysis at Full Vocabulary
After assessing the effectiveness of prefilling, we bring more insights into the behavior of considered LLMs in an unconstrained decoding setting and we analyze what happens when we do not restrict scoring to a predefined set of symbolic answer tokens. Instead, we allow the model to generate freely from its entire vocabulary and record the top-1 predicted token. We then evaluate whether this token corresponds to a valid answer label (e.g., ‘A’, ‘B’, ‘C’, ‘D’), and whether it matches the correct answer for the given question. For this analysis, we report results using two metrics: the full-vocabulary accuracy, which captures how often the real top-1 token is both valid and correct – without any filtering, and the First-Token Validity Rate (FTVR), which measures the proportion of examples where the top-1 token is a valid symbolic option.
As shown in Table 2 and Table 7 of the Appendix, applying prefilling substantially improves both FTVR and accuracy across all models. Notably, models such as Gemma-2-9B and Llama-3.1-8B experience gains of over 60 points. These results confirm that prefilling effectively shifts the full next-token distribution toward structured symbolic output, making first-token decoding a more reliable proxy for answer selection in MCQA tasks.
@@
Table 2: Full-vocabulary first-token evaluation on MMLU, OBQA, SIQA, and SCIQ. The model must predict a valid and correct token as its very first output. Prefilling improves both validity rate and accuracy.
@@ 4.5 Calibration Analysis
To further motivate our prefilling strategy, we evaluate the calibration of model predictions, focusing on how well the model confidence scores align with their actual accuracy. A well-calibrated model is one whose predicted confidence closely reflects the true likelihood of correctness. For example, predictions made with 80% confidence should be correct roughly 80% of the time. When this relationship breaks down, it indicates miscalibration: overconfidence occurs when confidence exceeds accuracy, while underconfidence reflects the opposite. Improving calibration is critical for increasing the reliability of model outputs, particularly in high-stakes or trust-sensitive applications.
To quantify calibration, we use two standard metrics: Expected Calibration Error (ECE) and the Brier score. Specifically, ECE measures the average discrepancy between predicted confidence and actual accuracy across discrete confidence bins. Formally, it is defined as
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|,$ (3)
where $M$ is the number of bins, $B_{m}$ is the set of predictions falling into bin $m$, and $\text{acc}(B_{m})$ and $\text{conf}(B_{m})$ denote the empirical accuracy and average confidence in that bin, respectively. Lower values indicate better calibration.
MMLU OBQA SIQA SCIQ
ECE ($\downarrow$) ECE ($\downarrow$) ECE ($\downarrow$) ECE ($\downarrow$) $\bar{\Delta}$
Llama-3.1-8B 20.6 45.7 24.9 1.1
+ prefilling 13.0 44.4 14.2 0.7 -5.0
Qwen-2-7B 24.6 51.7 20.3 2.1
+ prefilling 21.3 47.2 17.9 1.7 -2.7
Mistral-Nemo-12B 17.3 45.9 15.9 2.3
      + prefilling 16.3 44.8 14.1 1.1 -1.3
Phi-4-14B 36.7 50.6 21.1 3.5
      + prefilling 36.4 45.8 17.8 1.5 -2.6
Table 3: ECE across four MCQA benchmarks, before and after applying prefilling. Prefilling consistently lowers the ECE, indicating improved alignment between model confidence and actual correctness.
MMLU OBQA SIQA SCIQ
Brier-S ($\downarrow$) Brier-S ($\downarrow$) Brier-S ($\downarrow$) Brier-S ($\downarrow$) $\bar{\Delta}$
Llama-3.1-8B 23.4 47.5 28.4 2.2
+ prefilling 18.7 45.5 19.2 1.7 -4.1
Qwen-2-7B 25.2 50.8 21.8 2.1
+ prefilling 23.0 46.3 19.8 1.8 -2.25
Mistral-Nemo-12B 21.3 45.9 20.6 2.3
+ prefilling 20.4 46.3 18.5 1.9 -0.75
Phi-4-14B 36.4 50.3 24.1 3.7
+ prefilling 36.6 47.4 19.6 1.4 -2.38
Table 4: Brier score across four MCQA benchmarks, before and after applying prefilling. Prefilling consistently lowers the Brier score, indicating improved alignment between model confidence and actual correctness.
The Brier score, on the other hand, captures the mean squared difference between predicted probabilities and binary outcomes. It is given by
Brier $\text{Brier Score}=\frac{1}{n}\sum_{i=1}^{n}(p_{i}-y_{i})^{2},$ (4)
where $p_{i}$ is the model confidence in the correct answer and $y_{i}$ is the true label. A lower Brier score indicates both accurate and well-calibrated predictions, while a higher score penalizes incorrect predictions made with high certainty.
@@
We report the ECE scores in Table 3 and present the calibration diagrams in Figure 4. The former shows that prefilling consistently improves model calibration. For instance, Llama-3.1-8B exhibits notably reduced calibration error on datasets like MMLU and SocialIQA, with ECE dropping from 20.6% to 13.0% and from 24.9% to 14.2%, respectively. The diagrams in Figure 4 provide a more detailed view across different confidence levels. We observe that models such as Gemma-2-9B and Zephyr-7B tend to be overconfident in the mid-confidence range, while Llama-3.1-8B remains overconfident even at maximum predicted confidence (100%).
Brier scores are reported in Table 3. In all cases, prefilling reduces the Brier score, confirming that models become not only more accurate but also less overconfident in their incorrect predictions. This is especially noticeable on SIQA, where overconfident errors are common: for instance, Llama-3.1-8B improves from 28.4 to 19.2, and Phi-4-14B from 24.1 to 19.6. The average score reduction ($\bar{\Delta}$) across tasks highlights that the improvement is consistent across both mid- and high-capacity models.
This calibration analysis suggests that, in addition to improving accuracy, prefilling also enhances the trustworthiness of model confidence estimates.
@@
Figure 4: Calibration curves comparing standard FTP and FTP with prefilling. Prefilling improves calibration, bringing predictions closer to the ideal confidence – accuracy alignment.

### v2

@@ 4 Experiments
4.1 Experimental Setup
Benchmarks. We evaluate our approach across several widely recognized MCQA benchmarks, each designed to test different cognitive capabilities and domains of knowledge. We begin with general knowledge, assessed using MMLU [13], a comprehensive benchmark covering 57 academic and professional subjects. For reading comprehension, we employ RACE [18], a dataset derived from middle and high school exams and OpenBookQA (OBQA) [25], which requires reasoning over scientific facts and general knowledge. In the domain of commonsense reasoning, we include Social IQa (SIQA) [36] for social interaction understanding, Moral Stories (MS) [10] for moral decision-making and CommonsenseQA (CQA) [37] for everyday commonsense reasoning. For narrative understanding, we consider Story Cloze (SC) [28], where the model selects the most plausible story continuation, HellaSwag (HS) [47], which evaluates narrative coherence in the presence of adversarial distractors, and MC-TACO (MC-T) [50], which assesses temporal and causal reasoning in short texts. In the area of STEM, we use SciQ [44] for scientific knowledge and MathQA (MQA) [2] to evaluate mathematical problem-solving. For logical and analytical reasoning, we consider AI2 Reasoning Challenge (ARC) [8], a large-scale benchmark of science exam questions. We use both the Easy partition (ARC-E), which mainly involves fact recall and can often be answered through direct retrieval, and the Challenge partition (ARC-C), which requires multi-step reasoning and the integration of scientific knowledge. We also include LogiQA (LQA) [22], based on national civil service exams. Together, these datasets enable a comprehensive evaluation of our method effectiveness and robustness across varied cognitive demands.
Models. We assess a range of open-source instruction-tuned LLMs, selected to cover diverse architectures, sizes, and alignment strategies. Our pool includes Meta’s Llama-3.1-8B, Alibaba’s Qwen-2-7B, DeepMind’s Gemma-7B and Gemma-2-9B, HuggingFace’s DPO-tuned Zephyr-7B, Microsoft’s efficient Phi-4-14B, and two Mistral-family models: Ministral-8B and the alignment-focused Mistral-Nemo-12B. This selection enables a comprehensive analysis of prefilling effectiveness across different instruction-tuned models.
Prompt Format. For each benchmark, inputs are formatted as a concatenation of general instruction, question, and answers. To apply our prefilling strategy, we insert a natural-language prefix at the beginning of the model response. The default template used is: ‘Given the question and the possible options, my answer is: ’, placed immediately after the assistant message tag, which promotes symbolic alignment and structured outputs. This template was selected based on preliminary trials, as it consistently performed well across benchmarks and models. We consider as valid answers only token sequences that exactly match the symbolic labels (e.g., ‘A’, ‘B’, ‘C’, ‘D’), optionally preceded by up to two whitespace or newline characters. An analysis of robustness to alternative templates is provided in the supplementary material.
Table 1: Benchmarking base LLMs on natural language understanding, question answering, and reasoning tasks. Prefilling each model response positively impacts its ability to output the correct option, also against a direct prompt-instruction baseline. $\bar{\Delta}$ values denote average gains over base LLMs.
@@ 4.1 Experimental Setup
MMLU RACE OBQA SIQA MS CQA SC HS MC-T SciQ MQA ARC-E ARC-C LQA $\bar{\Delta}$
Llama-3.1-8B 63.1 78.1 71.2 70.9 87.0 72.7 93.2 52.2 82.1 97.1 24.0 90.2 52.7 28.9
+ prompting 63.7 78.2 80.4 71.9 90.1 77.2 94.8 49.2 89.7 97.8 24.0 78.7 49.8 28.4
+ prefilling 68.4 83.2 84.8 72.3 93.2 76.5 96.4 69.0 82.5 98.3 33.8 94.7 60.2 34.0 +6.0
Qwen-2-7B 68.9 87.0 83.8 75.6 88.1 79.4 97.1 53.4 88.3 97.5 29.8 94.3 48.0 33.0
+ prompting 68.3 86.7 84.2 75.2 85.3 79.4 97.7 65.0 85.1 97.2 29.8 74.1 47.9 31.6
+ prefilling 69.1 86.7 84.8 76.1 92.6 79.4 97.3 65.2 86.1 97.4 36.5 94.7 56.6 35.5 +2.4
Gemma-7B 45.9 50.4 40.6 34.2 77.5 51.6 69.6 36.3 60.1 89.0 24.1 55.9 43.9 26.4
+ prompting 48.1 65.2 63.8 57.5 65.7 64.7 77.2 43.9 91.0 93.6 24.0 58.7 37.3 27.3
+ prefilling 52.4 70.5 71.2 65.2 64.3 68.6 92.0 56.3 91.0 95.1 25.1 86.7 49.1 26.1 +14.9
Gemma-2-9B 33.9 49.5 38.8 56.3 85.9 39.6 97.6 43.0 72.2 31.8 27.8 38.6 52.4 31.0
+ prompting 70.2 86.2 87.6 74.2 91.6 78.7 97.9 63.6 77.1 98.1 27.8 81.2 55.9 32.1
+ prefilling 72.1 86.3 89.8 74.2 87.3 79.0 97.9 69.8 78.5 98.3 34.0 96.6 60.8 35.8 +25.9
Zephyr-7B 57.5 86.5 73.0 67.5 83.6 66.0 96.0 35.0 69.1 88.5 23.3 95.6 51.0 32.1
+ prompting 55.9 72.3 67.6 66.5 85.3 54.4 91.7 40.0 85.2 89.6 23.3 65.7 41.8 30.1
+ prefilling 58.8 87.3 70.8 68.5 86.2 69.0 96.0 36.1 93.3 93.1 24.3 98.1 51.0 40.6 +3.4
Ministral-8B 62.1 84.7 82.4 74.7 87.3 74.3 97.4 81.1 91.5 97.4 23.1 93.6 48.1 28.4
+ prompting 57.0 83.8 76.4 72.3 85.7 69.6 94.0 68.6 90.0 97.3 23.1 76.1 47.6 29.0
+ prefilling 63.9 84.7 85.2 75.9 89.9 75.0 98.2 86.5 91.9 98.0 24.5 93.7 49.1 34.6 +1.8
Mistral-Nemo-12B 65.1 82.7 80.6 74.2 78.0 74.1 97.3 51.6 92.4 96.8 23.5 92.6 51.3 28.7
+ prompting 58.1 79.5 69.8 68.1 74.2 68.8 89.9 52.6 92.8 95.9 23.5 76.5 50.0 31.9
+ prefilling 66.0 83.1 80.8 75.8 80.9 75.4 96.9 76.4 93.3 97.2 25.4 93.1 54.7 32.0 +3.0
Phi-4-14B 76.4 73.3 84.0 71.7 83.4 72.5 98.2 61.2 82.5 95.4 25.0 87.8 46.4 31.9
+ prompting 78.8 77.5 91.4 76.0 93.1 79.8 98.7 85.6 82.0 98.2 24.9 73.1 48.6 34.1
+ prefilling 79.7 73.4 90.0 77.3 93.2 79.6 98.7 87.8 86.1 98.3 45.0 87.8 60.2 34.6 +7.3
@@
We begin by evaluating the overall effectiveness of our prefilling strategy in enhancing the answer accuracy of our tested models across a diverse set of MCQA benchmarks, by comparing it with two natural baselines, the first one of which is the standard FTP approach. Specifically, for standard FTP, the model is prompted with the task and options, and the next-token probability distribution is used to score the first-token likelihoods assigned to each answer option. No prefix is injected. When using our prefilling strategy, instead, a fixed natural-language prefix is injected at the beginning of each assistant response, before the model generates any output. The FTP scoring is then applied exactly as in the baseline. This approach retains the efficiency of FTP but adds a steering mechanism that helps the model align with the expected symbolic format, while adding no extra latency to the generation. In addition to the standard FTP, we include a prompt-engineering-based baseline, where the instruction ‘Please answer only with [OPTIONS LIST]’ is appended to the original prompt. This allows us to directly compare output-side prefilling with prompt-side guidance.
Results. Table 1 reports first-token accuracy across benchmarks, comparing base LLMs with prompt-side additions and with our output prefilling strategy. Prefilling yields consistent improvements across all models and datasets. The largest average gain is observed for Gemma-2-9B, which improves by +25.9 points compared to standard FTP. Substantial improvements are also obtained for Gemma-7B (+14.9) and Zephyr-7B (+3.4). Even already strong models, such as Llama-3.1-8B and Phi-4-14B, achieve meaningful gains (+6.0 and +7.3, respectively), highlighting that prefilling is beneficial regardless of baseline capability.
In contrast, prompt-side additions lead to only modest or inconsistent improvements. Prefilling consistently outperforms prompt engineering, providing larger and more reliable accuracy gains. For example, Llama-3.1-8B improves by +10.4 points on ARC-C (60.2 vs. 49.8), Gemma-7B by +5.3 on RACE (70.5 vs. 65.2), Mistral-Nemo-12B by +23.8 on HS (76.4 vs. 52.6), and Phi-4-14B by +20.1 on MQA (45.0 vs. 24.9). These results demonstrate that mechanically enforcing the output format through prefilling is both more robust and broadly more effective than prompt-side instructions. These results confirm that structured prefilling reliably aligns the most probable model output with the intended symbolic answer, thereby reducing both misalignment and misinterpretation errors.
@@ 4.3 Alignment to Open-Ended Generation
Figure 3: Comparison of model accuracy on OpenBookQA, Social IQa, and SciQ using standard FTP, FTP with prefilling, and open-ended generation with different classifiers. FTP with prefilling consistently outperforms standard FTP and often surpasses more expensive open-ended approaches.
To further contextualize our results, we compare our approach against an open-ended generation setting, which we consider a gold-standard reference due to its flexibility and alignment with natural model behavior – albeit at higher computational, training, and evaluation cost due to decoding and external LLM classifier needs. In this setting, models generate free-form answers to each question without being constrained to select from symbolic options.
These open-ended responses are then mapped to symbolic answer labels through automatic evaluation using language model classifiers, such as GPT-3.5-Turbo [32], Llama-3.1-70B-Instruct [9], and the xFinder-Qwen classifier [46], which is specifically designed and fine-tuned for robust and precise answer extraction from LLM outputs. The aim of this comparison is to measure how well our prefilling strategy approximates the answers that a model would naturally produce in an unconstrained, expressive scenario. As in previous settings, we use accuracy to quantify model performance by evaluating whether the extracted symbolic answer matches the ground-truth.
Results. Fig. 3 reports results on three representative benchmarks: OpenBookQA, Social IQa, and SciQ [^1] The complete results on all benchmarks are reported in the supplementary material, as well as the exact prompting and classification details for open-ended responses.. We find that FTP with our prefilling strategy consistently improves alignment with symbolic labels derived from open-ended generation, outperforming the standard FTP baseline across all tasks. In several cases, it even surpasses open-ended generation under GPT- and xFinder-based evaluations, indicating a more accurate and efficient surrogate for expressive generation.
These results reinforce the idea that prefilling bridges the gap between efficient token-level scoring and natural free-form output, enabling structured symbolic accuracy without sacrificing alignment to the model unconstrained reasoning.
@@ 4.4 Analysis at Full Vocabulary
After assessing the effectiveness of prefilling, we bring more insights into the behavior of considered LLMs in an unconstrained decoding setting and we analyze
Table 2: Full-vocabulary first-token evaluation on MMLU, OBQA, SIQA, and SCIQ. The model must predict a valid and correct token as the first output. Prefilling improves both validity rate and accuracy.
@@
what happens when we do not restrict scoring to a predefined set of symbolic answer tokens. Instead, we allow the model to generate freely from its entire vocabulary and record the top-1 predicted token. We then evaluate whether this token corresponds to a valid answer label (e.g., ‘A’, ‘B’, ‘C’, ‘D’), and whether it matches the correct answer for the given question. For this analysis, we report results using two metrics: the full-vocabulary accuracy, which captures how often the real top-1 token is both valid and correct without any filtering, and the First-Token Validity Rate (FTVR), which measures the proportion of examples where the top-1 token is a valid symbolic option.
As shown in Table 2 and Table 6 of the supplementary material, applying prefilling substantially improves both FTVR and accuracy across all models. Notably, models such as Gemma-2-9B and Llama-3.1-8B experience gains of over 60 points. These results confirm that prefilling shifts the full next-token distribution toward structured symbolic output, making first-token decoding a more reliable proxy for answer selection in MCQA tasks.
Table 3: Calibration results on four MCQA benchmarks in terms of Adaptive Calibration Error (ACE), Brier score, and Log Loss, before and after applying prefilling. Prefilling consistently achieves lower values across all metrics, indicating improved alignment between model confidence and actual correctness. For all metrics, lower the better ($\downarrow$).
MMLU OBQA SIQA SCIQ
ACE Brier-S LogLoss ACE Brier-S LogLoss ACE Brier-S LogLoss ACE Brier-S LogLoss
Llama-3.1-8B 0.206 23.4 0.815 0.444 47.5 1.987 0.169 28.4 0.795 0.007 2.2 0.077
+ prefilling 0.129 18.7 0.616 0.252 45.5 1.341 0.138 19.2 0.752 0.006 1.7 0.048
Qwen-2-7B 0.246 25.2 1.311 0.327 50.8 1.790 0.210 21.8 1.213 0.020 2.1 0.112
+ prefilling 0.213 23.0 0.894 0.289 46.3 1.495 0.190 19.8 1.126 0.014 1.8 0.096
Mistral-Nemo-12B 0.173 21.3 0.673 0.288 45.9 1.414 0.157 20.6 0.707 0.007 2.3 0.079
+ prefilling 0.164 20.4 0.624 0.252 46.3 1.358 0.115 18.5 0.600 0.004 1.9 0.064
Phi-4-14B 0.187 36.4 1.084 0.340 50.3 1.712 0.210 24.1 1.547 0.034 3.7 0.177
+ prefilling 0.167 36.6 1.044 0.272 47.4 1.421 0.191 19.6 1.438 0.013 1.4 0.088
@@ 4.5 Calibration Analysis
To further motivate our prefilling strategy, we evaluate prediction calibration – i.e., how well confidence scores align with actual accuracy. A model is well-calibrated if predicted confidence reflects the true likelihood of correctness (e.g., confidence 80% implies $\sim$80% accuracy). When this relationship breaks down, it indicates miscalibration: overconfidence occurs when confidence exceeds accuracy, while underconfidence reflects the opposite. Improving calibration is critical for increasing the reliability of model outputs, particularly in high-stakes or trust-sensitive applications.
To quantify calibration, we use three metrics: Adaptive Calibration Error (ACE) [31], Brier score (Brier-S) [11], and Log Loss [12, 19]. ACE improves upon traditional metrics like Expected Calibration Error (ECE) [29] by addressing issues related to fixed binning and multi-class settings. Lower ACE values indicate better calibration. The Brier-S, on the other hand, captures the mean squared difference between predicted probabilities and binary outcomes. A lower Brier-S indicates both accurate and well-calibrated predictions, while a higher score penalizes incorrect predictions made with high certainty. Log Loss, instead, measures the discrepancy between predicted probabilities and true labels across all classes. It is particularly sensitive to overconfident incorrect predictions, complementing the Brier-S by emphasizing mistakes made with high certainty. The formal definitions of each metric are provided in the supplementary material.
The results are reported in Table 3, showing that prefilling consistently improves model calibration. In particular, examining the ACE metric, Llama-3.1-8B improves substantially on MMLU (from 0.206 to 0.129) and OpenBookQA (from 0.444 to 0.252). Even stronger baselines such as Phi-4-14B benefit from prefilling, with ACE on OpenBookQA dropping from 0.340 to 0.272. Consistent gains are also observed in terms of Brier-S, confirming that models become not only more accurate but also less overconfident in their incorrect predictions. This is especially noticeable on Social IQa, where overconfident errors are common: for instance, Llama-3.1-8B improves from 28.4 to 19.2, and Phi-4-14B from 24.1 to 19.6. Turning to Log Loss, prefilling reduces scores across all benchmarks and models. For example, on OpenBookQA, Llama-3.1-8B decreases from 1.987 to 1.341 while Phi-4-14B from 1.712 to 1.421. Since Log Loss heavily penalizes overconfident errors, these reductions indicate that prefilling helps models produce probability estimates that are better calibrated and more reliable.
@@
The calibration diagrams in Fig. 4 provide a more detailed view across different confidence levels. We observe that models such as Gemma-2-9B and Zephyr-7B tend to be overconfident in the mid-confidence range, while Llama-3.1-8B remains overconfident even at maximum predicted confidence (100%). This calibration analysis suggests that, in addition to improving accuracy, prefilling also enhances the trustworthiness of model confidence estimates.
@@
Figure 4: Calibration curves comparing standard FTP and FTP with prefilling. Prefilling improves calibration, moving predictions closer to the ideal confidence.